# Task Arithmetic Through The Lens Of One-Shot Federated Learning

### v1

@@ Abstract
Task Arithmetic is a model merging technique that enables the combination of multiple models’ capabilities into a single model through simple arithmetic in the weight space, without the need for additional fine-tuning or access to the original training data. However, the factors that determine the success of Task Arithmetic remain unclear. In this paper, we examine Task Arithmetic for multi-task learning by framing it as a one-shot Federated Learning problem. We demonstrate that Task Arithmetic is mathematically equivalent to the commonly used algorithm in Federated Learning, called Federated Averaging (FedAvg). By leveraging well-established theoretical results from FedAvg, we identify two key factors that impact the performance of Task Arithmetic: data heterogeneity and training heterogeneity. To mitigate these challenges, we adapt several algorithms from Federated Learning to improve the effectiveness of Task Arithmetic. Our experiments demonstrate that applying these algorithms can often significantly boost performance of the merged model compared to the original Task Arithmetic approach. This work bridges Task Arithmetic and Federated Learning, offering new theoretical perspectives on Task Arithmetic and improved practical methodologies for model merging.
@@ 1 Introduction
With the proliferation of fine-tuned models across diverse domains, efficiently combining these models to achieve excellence across multiple tasks has emerged as a critical research challenge. Task Arithmetic [28], a recent technique in model merging, offers a simple yet effective solution. For each fine-tuned model, a task vector is generated by subtracting the pre-trained model parameters from the fine-tuned model parameters. Summing these task vectors produces a direction that enhances performance of the pre-trained model across multiple tasks for which the fine-tuned models have been trained. A key advantage of this approach is that it only involves element-wise operations in the weight space, eliminating the need for additional fine-tuning.
@@
Despite its strong empirical performance, Task Arithmetic lacks substantial theoretical understanding. Only a small number of works have investigated this empirical success theoretically [55, 57]. In this paper, we take a step towards bridging the gap between theory and practice by framing Task Arithmetic as a form of one-shot Federated Learning.
@@
Federated Learning [52], a distributed machine learning paradigm, enables devices to collaboratively train one shared model without exchanging the raw data. Federated Learning goal is to retain data privacy and reduce computational costs, as all raw data remains stored locally on edge devices. In a typical Federated Learning training process, a server coordinates the training process by iterating through the following steps [32]. First, the server broadcasts the current global model parameters and a training program to all the devices. Then each device locally computes an update to the model by using its own data. Finally, the server aggregates all the local updates from devices and updates the current global model by using the aggregated local updates. A commonly used algorithm for this training process is Federated Averaging (FedAvg) [52]. In one-shot Federated Learning, the server learns a global model in only a single round of communication between itself and all the devices [23].
@@
We show that using one-shot FedAvg is equivalent to Task Arithmetic, thus offering a new perspective on Task Arithmetic through the lens of one-shot Federated Learning. Using the connection between Federated Learning and Task Arithmetic, we can leverage the extensive theoretical and algorithmic advancements in Federated Learning to better understand when Task Arithmetic is effective and how it can be improved. To the best of our knowledge, this is the first study to bridge Federated Learning and Task Arithmetic. Our main contributions are summarized as follows.
@@
  • Bridge Task Arithmetic and Federated Learning: We establish the connection between Task Arithmetic and one-shot Federated Averaging, formalizing Task Arithmetic using notions from Federated Learning.
@@
  • Analyze the Impact of Data and Training Heterogeneity in Task Arithmetic: Data heterogeneity slows convergence in FedAvg, while training heterogeneity causes objective inconsistencies. We show that similar challenges exist in Task Arithmetic and analyze their impact using insights from Federated Learning, offering a deeper understanding of its convergence behavior.
  • Identify and Adapt Federated Learning Algorithms for Task Arithmetic: We identify and recommend Federated Learning algorithms to address heterogeneity challenges and enhance Task Arithmetic for better model merging performance.
@@
  • Experiments Show That Federated Learning Algorithms Often Improve Task Arithmetic: Experiments confirm that adapting Federated Learning algorithms often improves the merged model’s performance compared to Task Arithmetic.
@@ 2 Task Arithmetic is One-Shot FedAvg
To deepen our understanding of the mechanism behind Task Arithmetic in multi-task learning, we establish a connection in this section between one-shot FedAvg and Task Arithmetic.
Given $T$ tasks, the objective in multi-task learning is to train a model parameterized by $\theta$ that performs well across all $T$ tasks. This can be formulated as minimizing the following multi-task objective function:
@@
Here, $\theta_{r}$ represents the parameter of the global objective function at the end of the $r$-th communication round, while $\theta_{t,r}^{(k)}$ denotes the parameter of the $t$-th local objective function at the $k$-th local optimization step during the $r$-th communication round. The learning rate used for this step is $\eta_{t,r}^{(k)}$. The stochastic gradient of the $t$-th local objective function $L_{t}$ is $g_{t}(\cdot)$, and $\beta$ is the outer step size used to aggregate all local updates. In the one-shot setting where $R=1$, the update simplifies to the following:
@@
where $\theta_{OS}$ denotes the parameters generated by one-shot Federated Learning.
In Task Arithmetic, the procedure mirrors the process in FedAvg. Each task $t$ independently minimizes its own objective function $L_{t}$ by performing $K_{t}$ iterations of SGD with learning rates $\{\eta_{t}^{(0)},\dots,\eta_{t}^{(K_{t}-1)}\}$, starting from the same initial model parameters $\theta_{0}$ and converging to a minimizer $\theta^{*}_{t}\in\arg\min L_{t}(\theta)$. This yields
$\theta^{*}_{t}=\theta_{t}^{(0)}-\sum_{k=0}^{K_{t}-1}\eta_{t}^{(k)}g_{t}(\theta_{t}^{(k)})$
where $\theta_{t}^{(0)}=\theta_{0}\,\forall t$. The task vector $\tau_{t}$ is defined by
$\tau_{t}=\theta_{t}^{*}-\theta_{t}^{(0)}=-\sum_{k=0}^{K_{t}-1}\eta_{t}^{(k)}g_{t}(\theta_{t}^{(k)}).$ (4)
Using Task Arithmetic, a new set of parameters can be constructed as
@@
where $\lambda$ is a hyperparameter, known as the scaling coefficient [28], which controls the extent to which the sum of task vectors is added back to the pre-trained parameters.
By comparing equations (3) and (5), we see that performing Task Arithmetic is equivalent to one-shot FedAvg with outer step size $\beta=\lambda T$.
@@ 3 Adapting Federated Learning Theory for Task Arithmetic
In this section, we extend theoretical insights from Federated Learning to Task Arithmetic, identifying the two main factors that impact its performance: data heterogeneity and training heterogeneity. Specifically, we analyze how these factors impact the convergence of Task Arithmetic. In particular, we study its ability to achieve the global minimum of a convex objective function and the local minimum of a non-convex objective function.
@@ 3.1 Data Heterogeneity
This subsection is dedicated to understanding how data heterogeneity influences the performance of Task Arithmetic. Data heterogeneity is common in Federated Learning and refers to the situation when data on each device is non-independent and identically distributed (non-i.i.d.) [42, 83, 70]. In Task Arithmetic, this issue is also prevalent since the training data associated with each task often comes from different distributions. In the convergence analysis of FedAvg, data heterogeneity has been a longstanding issue. Given the connection between FedAvg and Task Arithmetic, in order to understand how data heterogeneity impacts Task Arithmetic, it is helpful to first review existing findings on data heterogeneity in FedAvg. We begin by introducing several standard assumptions commonly used in the literature [43, 34, 35, 36, 73, 71, 21, 56, 68, 72].
@@
(Convexity and Smoothness) Assume all the task objective functions $L_{t}$ are convex and $H$-smooth. That is, $\forall t\in[T]$ and $\forall\theta,\varphi\in\mathbb{R}^{d}$,
$L_{t}(\theta)\leq L_{t}(\varphi)+\langle\nabla L_{t}(\varphi),\theta-\varphi\rangle+\frac{H}{2}\|\theta-\varphi\|^{2}.$
@@
(Bounded Initialization Error) Assume $\forall\theta^{*}\in\arg\min_{\theta\in\mathbb{R}^{d}}L(\theta)$, $\exists B$ such that the initialization $\theta_{0}$ satisfies
$\|\theta_{0}-\theta^{*}\|\leq B.$
To facilitate the analysis, we assume that all task objective functions are optimized with the same number of iterations, denoted $K_{t}=K\,\forall t\in[T]$, and that they use constant learning rates $\eta_{t}^{k}=\eta\,\forall t\in[T]$ $\forall k\in[K]$. Additionally, we set the outer step size to $\beta=1$, reducing Task Arithmetic to model averaging.
Although there is no universal definition of data heterogeneity, several notions are commonly referenced in the literature [43, 34, 35, 36, 73, 71, 21, 56, 72]. One widely adopted first-order notion of data heterogeneity is given by the following assumption [36, 56, 73, 21].
@@
The quantity $\zeta_{*}^{2}$ measures the diversity among the set of functions $\{L_{t}\}_{t=1}^{T}$ at the optima of the averaged multi-task objective function $L(\theta)$. Here, the notion of data heterogeneity is defined through objective functions, while [55] defines the Task Arithmetic property from the perspective of network functions. In Appendix A, we further explore the connection between this notion of data heterogeneity and the Task Arithmetic property proposed in [55].
Using the notation from [56], we define any learning problem that satisfies Assumptions 3.1, 3.2, 3.3, and 3.4 as belonging to the class $\mathcal{P}_{\zeta_{*}}^{H,B,\sigma}$. Building on the upper bound from [36] and the lower bound from [56], the following theorem characterizes the convergence rate of one-shot FedAvg.
@@
Here, $\succeq$ and $\preceq$ denote inequalities that hold up to absolute constants. Based on the above theorem, we make several observations about Task Arithmetic. First, data heterogeneity $\zeta_{*}^{2}$ degrades the performance of Task Arithmetic. The term $(H\zeta_{*}^{2}B^{4})^{1/3}$ is a non-vanishing error term introduced by $\zeta_{*}^{2}$, highlighting the impact of data heterogeneity.
Second, the one-shot learning nature of Task Arithmetic presents challenges that limit its performance. Notably, another non-vanishing term in Theorem 3.5, $HB^{2}$, arises due to the one-shot learning setup. In contrast, for FedAvg with $R$ communication rounds, this term becomes $\frac{HB^{2}}{R}$ and diminishes as the number of communication rounds $R$ increases. Moreover, although both $\frac{(H\sigma^{2}B^{4})^{1/3}}{K^{1/3}}$ and $\frac{\sigma B}{\sqrt{TK}}$ decrease as the number of local steps $K$ grows, they decay much more slowly in the one-shot setting compared to $R$ rounds of FedAvg. With multiple communication rounds, these terms are given by $\frac{(H\sigma^{2}B^{4})^{1/3}}{K^{1/3}R^{2/3}}$ and $\frac{\sigma B}{\sqrt{TKR}}$ respectively [56]. This underscores the additional challenges introduced by the one-shot learning paradigm.
Third, starting with a good pre-trained model is important. The influence of pre-training is captured by the term $B$ introduced in Assumption 3.3. This quantity $B$ is a critical factor as it appears in every error term, particularly in the non-vanishing term $(H\zeta_{*}^{2}B^{4})^{1/3}$. Starting with a well-suited pre-trained model that has a smaller $B$ significantly mitigates the adverse effects of high data heterogeneity $\zeta_{*}^{2}$, as a smaller $B$ counteracts the heterogeneity. In fact, the significance of pre-trained models has been observed in experiments of both Task Arithmetic [55] and Federated Learning [54, 6].
@@
Importance of scaling coefficient: As mentioned before, $\lambda=\frac{\beta}{T}$, meaning that the scaling coefficient depends directly on the outer step size. Although in this section we assume $\beta=1$ which yields $\lambda=\frac{1}{T}$ and reduces Task Arithmetic to model averaging for simplicity, actually proper tuning of the scaling coefficient $\lambda$ is essential. Research indicates that the choice of $\beta$ has a significant impact on FedAvg performance [56, 34, 4, 30, 49, 46]. A similar sensitivity to $\lambda$ has been observed in Task Arithmetic: the performance of the final model depends heavily on selecting the right $\lambda$. For instance, Figure 15 in [28] illustrates how Task Arithmetic’s performance can vary dramatically with changes to $\lambda$.

### v2

@@ Abstract
Task Arithmetic is a model merging technique that enables the combination of multiple models’ capabilities into a single model through simple arithmetic in the parameter space, without the need for additional fine-tuning or access to the original training data. However, the factors that determine the success of Task Arithmetic remain unclear. In this paper, we examine Task Arithmetic for multi-task learning by framing it as a one-shot Federated Learning problem. We demonstrate that Task Arithmetic is mathematically equivalent to the commonly used algorithm in Federated Learning, called Federated Averaging (FedAvg). By leveraging well-established theoretical results from FedAvg, we identify two key factors that impact the performance of Task Arithmetic: data heterogeneity and training heterogeneity. To mitigate these challenges, we adapt several algorithms from Federated Learning to improve the effectiveness of Task Arithmetic. Our experiments demonstrate that applying these algorithms can often significantly boost performance of the merged model compared to the original Task Arithmetic approach. This work bridges Task Arithmetic and Federated Learning, offering new theoretical perspectives on Task Arithmetic and improved practical methodologies for model merging. Our code is available at: https://github.com/SilviaTao/one_shot_fl_task_arithmetic
@@ 1 Introduction
With the proliferation of fine-tuned models across diverse domains, efficiently combining these models to achieve excellence across multiple tasks has emerged as a critical research challenge. Task Arithmetic (Ilharco et al., 2023), a recent technique in model merging, offers a simple yet effective solution. Consider a set of $T$ tasks, each defined by a training dataset and a loss function for fine-tuning. Let $\theta_{\operatorname{pre}}\in\mathbb{R}^{d}$ be the parameters of a pre-trained model and $\theta_{t}\in\mathbb{R}^{d}$ be the fine-tuned model parameters on task $t$. A task vector $\tau_{t}$ is generated by subtracting the pre-trained model parameters from the fine-tuned model parameters: $\tau_{t}=\theta_{\operatorname{pre}}-\theta_{t}$. The sum of these task vectors, $\sum_{t=1}^{T}\tau_{t}$, produces a direction that enhances performance of the pre-trained model across multiple tasks for which the fine-tuned models have been trained. In other words, updating the pre-trained model parameters by $\theta_{\operatorname{pre}}+\sum_{t=1}^{T}\tau_{t}$ results in a multi-task model. A key advantage of this approach is that it only involves element-wise operations in the weight space, eliminating the need for additional fine-tuning.
@@
Despite its strong empirical performance, Task Arithmetic lacks substantial theoretical understanding. Only a small number of works have investigated this empirical success theoretically (Ortiz-Jimenez et al., 2023; Porrello et al., 2024). In this paper, we take a step towards bridging the gap between theory and practice by framing Task Arithmetic as a form of one-shot Federated Learning.
@@
Federated Learning (McMahan et al., 2017), a distributed machine learning paradigm, enables devices to collaboratively train one shared model without exchanging the raw data. Federated Learning’s goal is to retain data privacy and reduce computational costs, as all raw data remains stored locally on edge devices. In a typical Federated Learning training process, a server coordinates the training process by iterating through the following steps (Kairouz et al., 2021). First, the server broadcasts the current global model parameters and a training program to all the devices. Then each device locally computes an update to the model by using its own data. Finally, the server aggregates all the local updates from devices and updates the current global model by using the aggregated local updates. A commonly used algorithm for this training process is Federated Averaging (FedAvg) (McMahan et al., 2017). In one-shot Federated Learning, the server learns a global model in only a single round of communication between itself and all the devices (Guha et al., 2019).
@@
In this work, we show that using one-shot FedAvg is equivalent to Task Arithmetic, thus offering a new perspective on Task Arithmetic through the lens of one-shot Federated Learning. Using the connection between Federated Learning and Task Arithmetic, we can leverage the extensive theoretical and algorithmic advancements in Federated Learning to better understand when Task Arithmetic is effective and how it can be improved. To the best of our knowledge, this is the first study to bridge Federated Learning and Task Arithmetic. Our main contributions and the structure of the paper are summarized as follows.
@@
Bridge Task Arithmetic and Federated Learning: In Section 3, we establish the connection between Task Arithmetic and one-shot Federated Averaging, formalizing Task Arithmetic by using notions from Federated Learning.
@@
Analyze the Impact of Data and Training Heterogeneity in Task Arithmetic: In Section 4, we use existing theoretical results from Federated Learning to deepen our understanding of Task Arithmetic. Specifically, we analyze how data heterogeneity, which slows convergence in FedAvg, also slows the convergence of Task Arithmetic (Section 4.1), and how training heterogeneity, which causes objective inconsistencies in Federated Learning, similarly impacts Task Arithmetic (Section 4.2).
Identify and Adapt Federated Learning Algorithms for Task Arithmetic: In Section 5, we adapt several Federated Learning algorithms to mitigate the challenges posed by data and training heterogeneity. To achieve effective adaptions, we first discuss challenges in adapting Federated Learning algorithms (Section 5.1), and then we recommend four Federated Learning algorithms to enhance Task Arithmetic (Section 5.2).
@@
Experimentally Show That Federated Learning Algorithms Often Improve Task Arithmetic: We conduct experiments on vision-language model CLIP (Radford et al., 2021) for image classification tasks in Section 6, and on large language model Llama2 (Touvron et al., 2023) for instruction following, mathematical reasoning and code generation in Section 7. Our experiments confirm that adapting Federated Learning algorithms often improves the merged model’s performance compared to Task Arithmetic.
2 Related Work
Our work bridges Federated Learning and Task Arithmetic, a prominent approach within the growing domain of model merging. Thus, this section reviews related work on both model merging and Federated Learning.
Model Merging Task Arithmetic is one of many recent works on model merging (Wortsman et al., 2022a; Frankle et al., 2020b; Wortsman et al., 2022b; Matena & Raffel, 2022; Ainsworth et al., 2022; Don-Yehiya et al., 2022). Though the term model merging is relatively new, firstly formalized by Matena & Raffel (2022), the concept has received significant investigation (Wortsman et al., 2022b; a; Izmailov et al., 2018; Yang et al., 2019; Tarvainen & Valpola, 2017; Gupta et al., 2020). For example, Wortsman et al. (2022b) average the pre-trained model parameters and fine-tuned parameters to enhance the robustness of fine-tuned model against distribution shifts. Wortsman et al. (2022a) show that averaging parameters of multiple fine-tuned models with different hyperparameter configurations can improve robustness and accuracy. Izmailov et al. (2018) demonstrate that averaging parameters along a single SGD trajectory leads to better generalization.
Task Arithmetic, introduced by Ilharco et al. (2023), refines model merging by introducing task vectors and a hyperparameter $\lambda$ to control how much task vectors modify pre-trained model parameters. This method has inspired various follow-up work on using simple arithmetic operations for model merging (Yu et al., 2024; Zhang et al., 2023; Ortiz-Jimenez et al., 2023; Stoica et al., 2023; Yadav et al., 2024a; Yang et al., 2023) such as sparsifying task vectors (Yu et al., 2024), merging parameter-efficient modules (Zhang et al., 2023), fine-tuning in linearized model spaces (Ortiz-Jimenez et al., 2023) and resolving sign interference of task vectors (Yadav et al., 2024a).
A concurrent work of Task Arithmetic is Jin et al. (2022), who propose the RegMean. When merging two linear regression models, the model merging problem can be formulated as an optimization problem which has a closed-form solution. RegMean uses this solution to merge every linear layer of fine-tuned models. Unlike Task Arithmetic, RegMean is limited to linear layers and requires access to the input features of all linear layers. While Jin et al. (2022) also note that model merging is an extreme case of Federated Learning with single communication round, they do not explore this further. Moreover, RegMean focuses on merging linear layers, distinguishing it from our framework which addresses more general model merging problems.
There is a substantial body of research dedicated to understanding the effectiveness of model merging (Draxler et al., 2018; Entezari et al., 2021; Frankle et al., 2020a; Garipov et al., 2018; Benton et al., 2021; Li et al., 2018; Foret et al., 2020; Simsek et al., 2021). Some studies focus on the theory of linear model connectivity (Draxler et al., 2018; Entezari et al., 2021; Frankle et al., 2020a; Garipov et al., 2018), while others emphasize the flatness of the loss landscape (Benton et al., 2021; Li et al., 2018; Foret et al., 2020; Simsek et al., 2021). However, there has been relatively little work addressing the effectiveness of Task Arithmetic except for Ortiz-Jimenez et al. (2023); Porrello et al. (2024). Our work aims to bridge this gap.
Federated Learning In Federated Learning, FedAvg (McMahan et al., 2017) is widely used to solve the following distributed optimization problem across $M$ devices: $\min_{\theta\in\mathbb{R}^{d}}L(\theta):=\frac{1}{M}\sum_{m=1}^{M}L_{m}(\theta)$ where $L_{m}(\theta):=\mathbb{E}_{x_{m}\sim\mathcal{D}_{m}}[\ell(\theta,x_{m})]$ is the objective function on each device $m$, defined by some loss function $\ell$ and data distribution $\mathcal{D}_{m}$. The core idea behind FedAvg is to perform local Stochastic Gradient Descent (local SGD) on each device, followed by model averaging on the server. A substantial body of research has analyzed the performance of FedAvg and local SGD (Li et al., 2019; Karimireddy et al., 2020; Khaled et al., 2020; Koloskova et al., 2020; Woodworth et al., 2020b; a; Glasgow et al., 2022; Patel et al., 2024; Wang et al., 2022; Woodworth et al., 2021).
A key challenge in the theoretical analysis of FedAvg arises from data heterogeneity, where each device $m$ has a different data distribution $\mathcal{D}_{m}$. In the homogeneous setting where $\mathcal{D}_{m}=\mathcal{D}\;\forall m$, Woodworth et al. (2021) have established the min-max complexity of FedAvg with smooth and convex loss functions. In the more complex heterogeneous setting, various works have derived the convergence rate of FedAvg under different assumptions about data heterogeneity (Li et al., 2019; Khaled et al., 2020; Koloskova et al., 2020; Woodworth et al., 2020b; a; Glasgow et al., 2022; Patel et al., 2024; Wang et al., 2022). In this work, rather than focusing on extending existing theoretical results, we focus on using these results to analyze Task Arithmetic.
To address the challenges posed by data heterogeneity, extensive research has focused on designing algorithms to improve the performance of FedAvg (Karimireddy et al., 2020; Reddi et al., 2020; Li et al., 2020b; Ye et al., 2023b; Li et al., 2021b; Tenison et al., 2022; Wang et al., 2020). Some work has enhanced optimization algorithms by regulating the local training process (Karimireddy et al., 2020; Reddi et al., 2020; Li et al., 2020b; Wang et al., 2020; Li et al., 2021b), while others have proposed alternative aggregation methods beyond simple averaging (Tenison et al., 2022; Ye et al., 2023b). Another line of work focuses on personalized Federated Learning (Smith et al., 2017; Tan et al., 2022; Mansour et al., 2020; Li et al., 2021a; Finn et al., 2017), addressing data heterogeneity by adapting the global model locally for each device.
Aside from data heterogeneity, Wang et al. (2020) notice different local training process (which we refer to training heterogeneity) exacerbates the convergence of federated optimization algorithms, leading them to converge to a stationary point of an objective function inconsistent with the original objective function.
@@ 3 Task Arithmetic is One-Shot FedAvg
To deepen our understanding of the mechanism behind Task Arithmetic in multi-task learning, in this section, we establish a connection between one-shot FedAvg and Task Arithmetic.
Given $T$ tasks, the objective in multi-task learning is to train a model parameterized by $\theta\in\mathbb{R}^{d}$ that performs well across all $T$ tasks. This can be formulated as minimizing the following multi-task objective function:
@@
Here, $\theta_{r}$ denotes the global model parameters obtained by optimizing the global objective at the end of the $r$-th communication round, and $\theta_{t,r}^{(k)}$ denotes the local model parameters obtained by optimizing the $t$-th local objective at the $k$-th local step during round $r$. The learning rate used for this step is $\eta_{t,r}^{(k)}$. The stochastic gradient of the $t$-th local objective function $L_{t}$ is $g_{t}(\cdot)$, and $\beta$ is the outer step size used to aggregate all local updates. In the one-shot setting where $R=1$, the update simplifies to the following:
@@
where $\theta_{OS}$ denotes the parameters generated by one-shot FedAvg.
In Task Arithmetic, the procedure mirrors the process in FedAvg. Each task $t$ independently minimizes its own objective function $L_{t}$ by performing $K_{t}$ iterations of SGD with learning rates $\{\eta_{t}^{(0)},\dots,\eta_{t}^{(K_{t}-1)}\}$, starting from the same pre-trained model parameters $\theta_{0}$ and converging to a minimizer $\theta^{*}_{t}\in\arg\min L_{t}(\theta)$. This yields $\theta^{*}_{t}=\theta_{t}^{(0)}-\sum_{k=0}^{K_{t}-1}\eta_{t}^{(k)}g_{t}(\theta_{t}^{(k)})$ where $\theta_{t}^{(0)}=\theta_{0}\,\forall t$. The task vector $\tau_{t}$ is defined by $\tau_{t}=\theta_{t}^{*}-\theta_{t}^{(0)}$. Using Task Arithmetic, a new set of parameters can be constructed as
@@
where $\lambda$ is a hyperparameter called the scaling coefficient (Ilharco et al., 2023), which controls the extent to which the sum of task vectors is added back to the pre-trained parameters. By comparing equations (3) and (4), we see that performing Task Arithmetic is equivalent to one-shot FedAvg with outer step size $\beta=\lambda T$.
@@ 4 Adapting Federated Learning Theory for Task Arithmetic
In this section, we extend theoretical insights from Federated Learning to Task Arithmetic, identifying two main factors that impact its performance: data heterogeneity and training heterogeneity. We analyze how these factors impact the convergence of Task Arithmetic to the global minimum of a convex objective function and to the local minimum of a non-convex objective function.
@@ 4.1 Data Heterogeneity
This subsection is dedicated to understanding how data heterogeneity influences the performance of Task Arithmetic. Data heterogeneity is common in Federated Learning and refers to the situation when data on each device is non-independent and identically distributed (non-i.i.d.) (Li et al., 2020b; Ye et al., 2023a; Wen et al., 2023). In the context of Task Arithmetic, we define data heterogeneity as follows:
Definition 4.1.
(Data Heterogeneity) In Task Arithmetic, data heterogeneity refers to the non-i.i.d nature of training data across different tasks.
In the convergence analysis of FedAvg, data heterogeneity has been a longstanding issue. Given the connection between FedAvg and Task Arithmetic, in order to understand how data heterogeneity impacts Task Arithmetic, it is helpful to first review existing findings on data heterogeneity in FedAvg. We begin by introducing several standard assumptions commonly used in the literature (Li et al., 2019; Karimireddy et al., 2020; Khaled et al., 2020; Koloskova et al., 2020; Woodworth et al., 2020b; a; Glasgow et al., 2022; Patel et al., 2024; Wang et al., 2022; Woodworth et al., 2021).
@@
(Convexity and Smoothness) Assume all the task objective functions $L_{t}$ are convex and $H$-smooth. That is, $\forall t\in[T]$ and $\forall\theta,\varphi\in\mathbb{R}^{d}$, $L_{t}(\theta)\leq L_{t}(\varphi)+\langle\nabla L_{t}(\varphi),\theta-\varphi\rangle+\frac{H}{2}\|\theta-\varphi\|^{2}.$
@@
(Bounded Initialization Error) Assume $\forall\theta^{*}\in\arg\min_{\theta\in\mathbb{R}^{d}}L(\theta)$, $\exists B$ such that the initialization $\theta_{0}$ satisfies $\|\theta_{0}-\theta^{*}\|\leq B$.
To facilitate the analysis, we assume that all task objective functions are optimized with the same number of iterations, denoted $K_{t}=K\,\forall t\in[T]$, and that they use constant learning rates $\eta_{t}^{k}=\eta\,\forall t\in[T]$ and $\forall k\in[K]$. Additionally, we set the outer step size to $\beta=1$, reducing Task Arithmetic to model averaging.
Although there is no universal definition of data heterogeneity, several notions are commonly referenced in the literature (Li et al., 2019; Karimireddy et al., 2020; Khaled et al., 2020; Koloskova et al., 2020; Woodworth et al., 2020b; a; Glasgow et al., 2022; Patel et al., 2024; Woodworth et al., 2021). One widely adopted first-order notion of data heterogeneity is given by the following assumption (Koloskova et al., 2020; Patel et al., 2024; Woodworth et al., 2020b; Glasgow et al., 2022).
@@
The quantity $\zeta_{*}^{2}$ measures the diversity among the set of functions $\{L_{t}\}_{t=1}^{T}$ at the optima of the averaged multi-task objective function $L(\theta)$. Here, the notion of data heterogeneity is defined through objective functions, while Ortiz-Jimenez et al. (2023) defines the Task Arithmetic property from the perspective of network functions. In Appendix A, we further explore the connection between this notion of data heterogeneity and the Task Arithmetic property proposed in Ortiz-Jimenez et al. (2023).
Using the notation from Patel et al. (2024), we define a learning problem satisfying Assumptions 4.2, 4.3, 4.4, and 4.5 as belonging to class $\mathcal{P}_{\zeta_{*}}^{H,B,\sigma}$. Based on the upper bound from Koloskova et al. (2020) and the lower bound from Patel et al. (2024), the following theorem characterizes the convergence of one-shot FedAvg.
@@
Here, $\succeq$ and $\preceq$ denote inequalities that hold up to absolute constants. Based on the above theorem, we make several observations about Task Arithmetic. First, data heterogeneity $\zeta_{*}^{2}$ degrades the performance of Task Arithmetic. The term $(H\zeta_{*}^{2}B^{4})^{1/3}$ is a non-vanishing error term introduced by $\zeta_{*}^{2}$, highlighting the negative impact of data heterogeneity.
Second, the one-shot learning nature of Task Arithmetic presents challenges that limit its performance. Notably, another non-vanishing term in Theorem 4.6, $HB^{2}$, arises due to the one-shot learning setup. In contrast, for FedAvg with $R$ communication rounds, this term becomes $\frac{HB^{2}}{R}$ and diminishes as the number of communication rounds $R$ increases. Moreover, although both $\frac{(H\sigma^{2}B^{4})^{1/3}}{K^{1/3}}$ and $\frac{\sigma B}{\sqrt{TK}}$ decrease as the number of local steps $K$ grows, they decay much more slowly in the one-shot setting compared to $R$ rounds of FedAvg. With multiple communication rounds, these terms are given by $\frac{(H\sigma^{2}B^{4})^{1/3}}{K^{1/3}R^{2/3}}$ and $\frac{\sigma B}{\sqrt{TKR}}$ respectively (Patel et al., 2024). This underscores the additional challenges introduced by the one-shot learning paradigm.
Third, starting with a good pre-trained model is important. The influence of pre-training is captured by the term $B$ introduced in Assumption 4.4. This quantity $B$ is a critical factor as it appears in every error term, particularly in the non-vanishing term $(H\zeta_{*}^{2}B^{4})^{1/3}$. Starting with a well-suited pre-trained model that has a smaller $B$ significantly mitigates the adverse effects of high data heterogeneity $\zeta_{*}^{2}$, as a smaller $B$ counteracts the heterogeneity. In fact, the significance of pre-trained models has been observed in experiments of both Task Arithmetic (Ortiz-Jimenez et al., 2023) and Federated Learning (Nguyen et al., 2022; Chen et al., 2022).
@@
The scaling coefficient $\lambda$ is important. As mentioned before, $\lambda=\frac{\beta}{T}$, meaning that the scaling coefficient depends directly on the outer step size. Although in this section we assume $\beta=1$ which yields $\lambda=\frac{1}{T}$ and reduces Task Arithmetic to model averaging for simplicity, proper tuning of the scaling coefficient $\lambda$ is essential. Research indicates that the choice of $\beta$ has a significant impact on FedAvg performance (Patel et al., 2024; Karimireddy et al., 2020; Charles & Konečnỳ, 2020; Jhunjhunwala et al., 2023; Malinovsky et al., 2023; Li et al., 2023b). A similar sensitivity to $\lambda$ has been observed in Task Arithmetic: the performance of the final model depends heavily on $\lambda$. For instance, Figure 15 in Ilharco et al. (2023) illustrates how Task Arithmetic’s performance can vary dramatically with changes to $\lambda$.